OpenStudy (anonymous):

Medal and a fan for help! (: What is the value of 2x−2y−2 for x=3 and y=-2?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

sorry, 8

OpenStudy (anonymous):

This can be represented by 2(3) - 2(-2) - 2 = 6 + 4 -2 = 8 Does that make sense?

OpenStudy (anonymous):

Well that's what I id at first but that not one of the answer choices. a. 1/18 b. -4 c. -1/32 d. -16

OpenStudy (anonymous):

\[2x ^{-2} y ^{-2}\]

OpenStudy (ybarrap):

$$ 2x^{-2}y^{-2}\\ 2(3)^{-2}(-2)^{-2}=(2/9)\times(1/4)=2/36=1/18\\ $$

OpenStudy (ybarrap):

make sense?

OpenStudy (anonymous):

Oh, okay yeah that's the answer I got first, I just needed to make sure it was right, okay ready for the next/last one? (:

OpenStudy (anonymous):

What is the value of \[\frac{ 1}{ 2^{-2}x ^{-3}y ^{5} }\] for x=2 and y= -4

OpenStudy (ybarrap):

$$ \large{ \frac{ 1}{ 2^{-2}x ^{-3}y ^{5} }\\ =\frac{ 1}{ 2^{-2}(2) ^{-3}(-4) ^{5} }\\ =\frac{ 1}{ (1/4)(1/8)1024 }\\ =\frac{ 1}{ (1/32)1024 }\\ =\frac{ 1}{ 32}\\ } $$

OpenStudy (ybarrap):

Any questions?

OpenStudy (ybarrap):

btw, you're not having problems viewing the equations? did you have to upgrade your browser or something?